OpenStudy (anonymous):

A rectangular garden is 7 yards wide. If you walked diagonally across the garden, you would walk 25 yards. How many yards long is the garden?

OpenStudy (anonymous):



OpenStudy (anonymous):

using Pythagoras theorem, can you find the length of the rectangle??

OpenStudy (anonymous):

25^2+b^2=7^2?

OpenStudy (anonymous):

625+b^2=49?

OpenStudy (anonymous):

it is rather \[25^{2} = 7^{2} + b ^{2}\]

OpenStudy (anonymous):

625=49+b^2 how do i get b ?

OpenStudy (anonymous):

is 24?

OpenStudy (anonymous):

yes